OpenStudy (anonymous):

find the discriminant of the following equation -12x^2-10x-11=0 and discribe the nature of the solutions one real solution two real solutions or, two imaginary solutions?

OpenStudy (watchmath):

\(D=b^2-4ac=(-10)^2-4(-12)(-11)=100-528=-428\) So the equatin have two imaginary solutions